OpenStudy (anonymous):

Calculus Help!! Find the sum of the series \[\sum_{n=0}^{\infty} 8^{n}/(9^{n}n!)\]

OpenStudy (anonymous):

i got 0 but it is incorrect!!

OpenStudy (perl):

use ratio test?

OpenStudy (perl):

first of all you have (8/9)^n

OpenStudy (anonymous):

yes

OpenStudy (experimentx):

it's a converging series ... but sum is obviously not zero.

OpenStudy (perl):

cant be zero, plug in a few terms LOL

OpenStudy (anonymous):

i got\[\lim_{n \rightarrow \infty}\] 8/9(n+1)

OpenStudy (perl):

no thats just a condition to make sure it does not diverge

OpenStudy (anonymous):

so it is not equal to sum?

OpenStudy (perl):

no, the limit is just the limit of the sequence terms. not adding them

OpenStudy (anonymous):

but how should i start with for adding them?

OpenStudy (perl):

suppose your sequence is {1/2^n} , n=1,2,3... so your sequence is 1/2, 1/4, 1/8, 1/16, ... lim (1/2^n) =0, does the series (the sum of the sequence terms) equal zero? of course not!!! 1/2 + 1/4 + 1/8 + ... clearly is not zero the question is, does it diverge ?

OpenStudy (anonymous):

it diverges!!

OpenStudy (perl):

does it? how do you know

OpenStudy (perl):

no it doesnt diverge, that is a geometric series , it converges

OpenStudy (anonymous):

i think the answer should be 9,am i right?

OpenStudy (anonymous):

no!!sorry!!

OpenStudy (anonymous):

is it should be a/(1-r)

OpenStudy (anonymous):

The answer is e^(8/9) e^x = Sum[ x^n/n! , {n,0, Infinity]] x = 8/9

OpenStudy (anonymous):

Thanks!!

OpenStudy (anonymous):

but is it you can only get it when you knew some of the patterns of the functions?

OpenStudy (experimentx):

somewhat seem to me like e times geometric progression of 8/9

OpenStudy (experimentx):

since it is less than 1 ... it is definitely going to be less than e

OpenStudy (anonymous):

it makes sense but in this case,we have to know e^x pattern in order to get this question?

OpenStudy (experimentx):

we know that 1+1+1/2+1/3!+1/4! = e +1

OpenStudy (anonymous):

okay!!

OpenStudy (experimentx):

also we know (8/9)^t is a geometric progression

OpenStudy (anonymous):

okay!!i think i got it!!Thanks everyone=))

OpenStudy (experimentx):

still we haven't calculated the value yet ... we have just deduced that it's less than e

OpenStudy (anonymous):

but is it you can only get it when you knew some of the patterns of the functions? Yes

OpenStudy (anonymous):

i put e^(8/9)...according to @eliassaab and it makes sense now.since the sum of the e^x is e^x/n!

OpenStudy (anonymous):

e^x = Sum[ x^n/n! , {n,0, Infinity]]

OpenStudy (experimentx):

yeah ... he's right

OpenStudy (perl):

ohh

OpenStudy (perl):

elia, is there a test you can use to see if this converges

OpenStudy (perl):

that was pretty cloever, i wouldnt have thought of that

OpenStudy (experimentx):

tests are lot easier than fiding sum, use comparison test ... for best

OpenStudy (anonymous):

@perl you can use the ratio test.

OpenStudy (perl):

right, this was a stronger result

OpenStudy (experimentx):

(8/9)^t/n! < 1/n! since 1/n! converges, the LHS must also converge

OpenStudy (anonymous):

you must know the patterns first!!

OpenStudy (perl):

but then , how do you prove series 1/n! converges?

OpenStudy (perl):

is that using ratio test?

OpenStudy (experimentx):

it's a standard value 1/n! = e

OpenStudy (perl):

true

OpenStudy (anonymous):

@experimentX you still need the ratio test to show that 1/n! converges

OpenStudy (anonymous):

it's a standard value 1/n! = e Sum[ 1/$n!, {n,0, Infinity] =e

OpenStudy (experimentx):

well, yes ... there it goes. 1/n!/1/(n+1)! -> 0 as n->inf

OpenStudy (perl):

so in other words, you can prove that series 1/n! converges. it turns out that it equals e

OpenStudy (anonymous):

If you use the ratio test with the original series, you do not need to do it again.

OpenStudy (experimentx):

yeah that's right too ...

OpenStudy (perl):

eliassab, i wanted to ask you about last question. it is sufficient to know that limit of f(x) exists at x =0. you dont need continuity there

OpenStudy (anonymous):

@perl yes.